OpenStudy (anonymous):

What is the solution to the equation? y – 0.7 = –2.2 y =

OpenStudy (anonymous):

2.5

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. y−0.7=−2.2 Step 1: Add 0.7 to both sides. y−0.7+0.7=−2.2+0.7 y=−1.5 Answer: y=−1.5 @fallengoddess